i have a really confusing question on my homework. it says find an equation of the line passing through the point (-5,2) with slope m=-4. if you could please help me out that would be great!
Answer by ankor@dixie-net.com(22740)

i have a really confusing question on my homework. it says find an equation of the line passing through the point (-5,2) with slope m=-4. if you could please help me out that would be great!
:
Just use the point/slope equation: y - y1 = m(x - x1)
:
In your problem: x1 = -5; y1 = 2, m = -4
:
y - 2 = -4(x - (-5)
:
y - 2 = -4(x + 5); remember a minus, a minus, is a plus
:
y -2 = -4x - 20
:
y = -4x - 20 + 2
:
y = -4x - 18 is the equation
:
:
:
Check this by substituting -5 for x and see if you get y = 2
y = -4(-5) - 18
y = +20 - 18